Capsule Review of “Comparing Id and Haskell in a Monte Carlo Photon Transport Code”

Jeffrey Hammes, Olaf Lubeck, Wim Boehm

The Monte Carlo technique has a long history. Its importance has grown in tandem with the availability of cheap computing power. The authors find that a particular large Monte Carlo simulation code can be naturally transcribed to a functional style. But there are caveats. It is characteristic of the Monte Carlo method that code validation and debugging depend on high-statistics results. The frank description of the problems involved in obtaining such results from the functional codes makes disturbing reading.